Hi, I needed some help for this, so if you could show me step by step with the answer, it'd be great.Thanks.
One grade of tea costing $3.20 per pound is mixed with another grade costing $2 per pound to make 20 lbs of a blend that will sell for $2.72 per pound. How much of the $3.20 is needed?

Answer by ptaylor(2198)

OK
Lex x=amount of $3.20 tea that is needed
then 20-x=amount of $2.00 tea that is needed
Now we know that the total cost of the tea before it is mixed together($3.20x+$2.00(20-x)) has to equal the total cost after it is mixed together($2.72*20), soooooo:
$3.20x+$2.00(20-x)=$2.72*20 simplify
$3.20x+$40.00-$2.00x=$54.40 subtract $40.00 from each side
$1.20x=$14.40
x=12 pounds---amount of $3.20 tea needed
20-x=20-12=8 pounds ----amount of $2.00 tea needed
Hope this helps---ptaylor